OpenStudy (anonymous):

A man borrows Rs 18000 at 5% per annum compound interest .If he repays Rs 6000 at the end of the first year and Rs8000 at the end of the second year; how much should he pay at the end of the 3rd year in order to clear the account?

OpenStudy (anonymous):

\[A_{1}=18000(1+0.05)-6000\] \[A_{2}=[18000(1.05)](1.05)-8000\] \[A_{3}=18000(1.05)^2-p\]

OpenStudy (anonymous):

Oops the second installment meant o have -6000 inside the brackets

OpenStudy (anonymous):

\[A_{2}=[18000(1.05)-6000](1.05)−8000\]

OpenStudy (anonymous):

A_{3}=[18000(1.05)^2-6000(1.05)-8000](1.05)-p.

OpenStudy (anonymous):

\[A_{3}=[18000(1.05)^2-6000(1.05)-8000](1.05)-p\]

OpenStudy (anonymous):

A1=18000(1+0.05)−6000 A2=[18000(1.05)−6000](1.05)−8000 A3=[18000(1.05)2−6000(1.05)−8000](1.05)−p

OpenStudy (anonymous):

Is the answer $5822.25?

OpenStudy (anonymous):

yup..how do u got it

OpenStudy (anonymous):

Find p. Don't worry about A3 that's just to tell you what year or installment it is.

OpenStudy (anonymous):

Just ignore the A3 and pretend it was zero. Now just find p.

OpenStudy (anonymous):

i didn't get ..if i ignore A3 then how would i find p

OpenStudy (anonymous):

\[[18000(1.05)^2−6000(1.05)−8000](1.05)=5822.25\] \[18000(1.05)^3-6000(1.05)^2-8000(1.05)=5822.25\] Just expand the brackets.

OpenStudy (anonymous):

Plug that into your calculator and you will get the answer.